OpenStudy (anonymous):

Solve the logarithm log x = 1 Solve the logarithm log x = 1 @Mathematics

OpenStudy (anonymous):

10

OpenStudy (anonymous):

Taking log(x) to the 10th power cancels out the log So raise ten to both sides of the equation 10^(logx)=10^1 x=10

OpenStudy (anonymous):

Oh so if there is no base, it is automatically 1?

OpenStudy (anonymous):

*10

OpenStudy (anonymous):

10